Average score of cricketer in 2 matches is 27 & in 3 others is 32. Then his average score in 5 matches is what?
Answer by mananth(16949)

Average score of cricketer in 2 matches is 27 & in 3 others is 32. Then his average score in 5 matches is what?
...
Average score of cricketer in 2 matches is 27
total score = 54
in 3 others is 32.
total score = 3*32=96
..
total score in all 5 matches = 96+54=150
number of matches = 5
average score = 150/5 = 30
Then his average score in 5 matches is 30
...
m.ananth@hotmail.ca